The area of a triangle is 35 square centimeters. the base is 3 centimeters greater than the height. which equation could you use to find the height of the triangle in centimeters?
Answer by Cromlix(4381)

Hi there,
Make height = x
Base = 3 + x
Area of triangle = 1/2(base x height)
35 = 1/2((3 + x)x)
35 = 1/2(3x + x^2)
Multiply both sides by 2
70 = 3x + x^
Rearrange in ax^2 + bx + c form
x^2 + 3x - 70 = 0
Factorize
(x + 10)(x - 7) = 0
x + 10 = 0
x = -10 (Disregard as -ve)
x - 7) = 0
x = 7
Base = 10 cm
Height = 7 cm.
Hope this helps :-)